# A Generative-First Neural Audio Autoencoder

### v1

@@ 1 Introduction
Latent generative models have revolutionized audio synthesis, enabling applications from music generation [5, 9, 16] to source separation [8], upmixing [3], and understanding tasks [10]. These models fundamentally depend on neural audio autoencoders to compress raw waveforms into tractable latent representations. However, existing autoencoders are primarily designed as quantized variational autoencoders for reconstruction tasks, then mildly adapted for generative modeling through post hoc modifications. This reconstruction-first design philosophy creates fundamental mismatches with generative requirements, leading to inefficient tokenization rates (we use “tokenize” to describe converting audio to any latent representation), fragmented architectures across audio channel formats, and computational bottlenecks that limit practical deployment at scale. The discrete-continuous latent representation divide further compounds these issues: discrete methods lack continuous latents for diffusion, and continuous methods lack discrete tokens for language models. These challenges are particularly acute for high-fidelity music processing at 44.1 kHz, where complex content amplifies the tension between compression, quality, and complexity.
These limitations are evident across established audio codecs. SoundStream, EnCodec, and DAC [23, 6, 12] exemplify this paradigm, operating at 75-150 Hz with only quantized latents unsuitable for diffusion training. A 4-minute song requires over 18,000 tokens, creating memory bottlenecks, while slow encoding can constitute 30% of training time, limiting data augmentation and throughput. Recent attempts achieve success with significant trade-offs: Stable Audio Open [9] reduces rates to 21.5 Hz but increases encoding costs and only provides continuous latents unsuitable for language model training, SpectroStream [14] maintains low rates but requires 64 codebooks, and HILCodec [1] improves speed but maintains high token rates. The Music2Latent [18] line of work, culminating in CoDiCodec [19], achieved impressive 11 Hz compression rates along with continuous and discrete latents, at the cost of performance in signal-level metrics. None of these explicitly account for different audio channel formats. Speech-focused approaches [4, 17, 7] achieve 12.5 Hz but target low-bandwidth applications unsuitable for high-fidelity music with its broader frequency content and complex stereo imaging. These developments point toward the need for unified architectures designed specifically for generative modeling.
@@ 2.3 Post-training optimization
Unified latent representations: To support both continuous and discrete latents, we use a two-stage procedure within a single model. First, we train a continuous latent representation (GenAE-KL). We then apply latent restructuring [3, 2] to learn an inner RVQ bottleneck, discretizing the representation without retraining the backbone (GenAE-VQ). While post-hoc adaptation in the reverse direction (VQKL) has been explored [24], we adopt the complementary KLVQ route for practical reasons: we and others have found that end-to-end VQ training at very low rates can be brittle and resource-intensive [17, 7], and using pre-quantizer latents from a discrete model can complicate diffusion training [24]. Adding RVQ post hoc leverages the stability of the continuous model and allows the same architecture to encode into and decode from either KL or VQ latents, providing flexibility for downstream applications and facilitating controlled comparisons between diffusion/flow models and language models.
@@ 3.3 Evaluation metrics
We measure processing speed via real-time factor (RTF) and reconstruction quality using scale-invariant signal-to-distortion-ratio (SI-SDR), multi-resolution short-time fourier-transform (STFT) loss, and mel-spectrogram L1 distance, all in bfloat16 precision. We choose bfloat16 over float32 or float64 as nearly all generative modeling is performed in reduced precision, and is our target application. Note that float32 yields significantly different values but similar trends. For baselines at non-44.1 kHz rates, we resample to native rates for processing and back to 44.1 kHz for evaluation. We use the Song Describer [15] dataset and filter out all vocals using a voice activity detection model. We only train on instrumental content and aim to only evaluate on instrumental content.
@@ 3.4 Baseline comparisons
We evaluate GenAE through speed, audio channel format, and latent format ablation studies. We compare with a suite of state-of-the-art codecs. Baselines include DAC, EnCodec (24/48 kHz), Stable Audio Open, and CoDiCodec (discrete/continuous).
@@ 4.1 Speed benchmark
Our generative-first design philosophy prioritizes encoding speed, recognizing that generative models require encoding massive datasets during training. This makes encoding speed key to training throughput and to on-the-fly data augmentations (e.g. stem remixing, pitch/time shifts) where precomputing all possible augmentation-combinations would be impractical. We evaluate design decisions using real-time factor (processed/elapsed time). Fig. 1 shows results for encoding and decoding 2 60-second stereo samples at 44.1 kHz. Starting from our DAC-like baseline with Snake activations (Base), we achieve progressive improvements through: efficient activations (+4.5% encode), pre-residual downsampling (+36.1% encode), separable convolutions (+6.5% encode), and aggressive temporal downsampling (13.9 encoding speedup). Additional components include mel-spectrogram fusion and windowed attention with format conditioning. The resulting GenAE model uses 3 less memory than SAO, encodes 12 faster than SAO, and decodes 1.6 faster than DAC, enabling faster processing for easier generative modeling.
@@ 4.2 Compression and quality benchmark
Effective generative modeling requires autoencoders that achieve aggressive compression and good reconstruction quality while supporting continuous and discrete latents, enabling diffusion or language modeling. Our evaluation demonstrates how GenAE achieves this joint objective. Fig. 3 shows GenAE achieving superior rate-distortion performance essential for generative modeling. At 13.125 Hz, GenAE surpasses SAO in all metrics while running at 60% of the rate, and matches DAC performance at just 15% of the rate. This directly enables faster training and reduced memory requirements for generative models. At 36.75 Hz, GenAE-KL outperforms all baselines across all metrics while maintaining substantial compression. GenAE-VQ at 13.125 Hz again outperforms SAO across all metrics, demonstrating unified architecture performance across both continuous and discrete representations. Note that CoDiCodec sounds better than the metrics show. This unification eliminates the need for separate models when comparing diffusion and language model approaches, addressing a benchmarking challenge [21]. We also evaluate PESQ-WB 16KHz as recommended in [13]. PESQ-WB follows mel/STFT/SI-SDR and keeps GenAE-KL on the rate-distortion frontier. GenAE-KL at 36.75 Hz is best overall (4.04), and GenAE-KL at 13.125 Hz is the high-compression point (3.00). Baseline scores: CoDiCodec (1.64), SAO (2.76), DAC (3.49), EnCodec-24 (3.67), EnCodec-48 (3.77).
@@ 4.3 Multi-format unification benchmark
Practical generative modeling requires autoencoders that work seamlessly across different audio channel configurations. Our unified architecture eliminates the need for multiple specialized models, reducing deployment complexity with consistent performance across major audio channel formats. Tab. 1 evaluates Mel reconstruction performance across L/R and M/S stereo audio channel formats. GenAE performs nearly identically in both -KL and -VQ forms across audio channel formats, demonstrating consistent quality regardless of input audio channel format or internal latent configuration. This is crucial for generative models to process diverse content. In contrast, CoDiCodec, SAO and EnCodec-48 degrade significantly due to L/R-specific training, while mono models are varied (DAC maintains, EnCodec-24 degrades). This highlights the importance of multi-format design since GenAE-based models can manipulate different audio formats without needing to re-encode/decode.
@@ 4.4 Generative context benchmark
| EnCodec-48 | 150 | 73 | 0.5485 | 0.6602 |
| DAC | 86 | 127 | 0.5144 | 0.5114 |
| CoDiCodec-FSQ | 11 | 993 | 0.9586 | 1.0553 |
@@ 5 Conclusion
In this work, we design neural audio autoencoders for generative models. Our GenAE is a single backbone that autoencodes at rates as low as 13.125 Hz ( downsampling), supports both continuous (KL) and discrete (VQ) latents, and operates across mono, stereo, and mid/side audio channel formats without model variants. A principled combination of efficient activations, early downsampling, separable convolutions, mel-spectrogram fusion, and windowed attention yields order-of-magnitude faster encoding and substantially lower memory while maintaining competitive fidelity, placing GenAE on the rate–distortion frontier with far fewer tokens. Our generative-first design repositions tokenization from a computational bottleneck to an efficient, scalable component: a 60-second track compresses to 788 tokens, enabling tractable long-context transformers. A post-training step lets us use both continuous and discrete latents in one model.

### v2

@@ 1 Introduction
Latent generative models have revolutionized audio synthesis, enabling applications from music generation [5, 9, 17] to source separation [8], upmixing [3], and understanding tasks [10]. These models fundamentally depend on neural audio autoencoders to compress raw waveforms into tractable latent representations. However, existing autoencoders are primarily designed as quantized variational autoencoders for reconstruction tasks, then mildly adapted for generative modeling through post-hoc modifications. This reconstruction-first design philosophy creates fundamental mismatches with generative requirements, leading to inefficient tokenization rates (we use “tokenize” to describe converting audio to any latent representation), fragmented architectures across audio channel formats, and computational bottlenecks that limit practical deployment at scale. The discrete-continuous latent representation divide further compounds these issues: discrete methods lack continuous latents for diffusion, and continuous methods lack discrete tokens for language models. These challenges are particularly acute for high-fidelity music processing at 44.1 kHz, where complex content amplifies the tension between compression, quality, and complexity.
These limitations are evident across established audio codecs. SoundStream, EnCodec, and DAC [26, 6, 13] exemplify this paradigm, operating at 75-150 Hz with only quantized latents unsuitable for diffusion training. A 4-minute song requires over 18,000 tokens, creating memory bottlenecks, while slow encoding can constitute 30% of training time, limiting data augmentation and throughput. Recent attempts achieve success with significant trade-offs: Stable Audio Open [9] reduces rates to 21.5 Hz but increases encoding costs and only provides continuous latents unsuitable for language model training, SpectroStream [15] maintains low rates but requires 64 codebooks, and HILCodec [1] improves speed but maintains high token rates. The Music2Latent [19] line of work, culminating in CoDiCodec [20], achieved impressive 11 Hz compression rates along with continuous and discrete latents, at the cost of performance in signal-level metrics. None of these explicitly account for different audio channel formats. Speech-focused approaches [4, 18, 7] achieve 12.5 Hz but target low-bandwidth applications unsuitable for high-fidelity music with its broader frequency content and complex stereo imaging. These developments point toward the need for unified architectures designed specifically for generative modeling. Concurrent works [25, 11] push music to 12.5 Hz on 24 kHz audio.
@@ 2.3 Post-training optimization
Unified latent representations: To support both continuous and discrete latents, we use a two-stage procedure within a single model. First, we train a continuous latent representation (GenAE-KL). We then apply latent restructuring [3, 2] to learn an inner RVQ bottleneck, discretizing the representation without retraining the backbone (GenAE-VQ). While post-hoc adaptation in the reverse direction (VQKL) has been explored [27], we adopt the complementary KLVQ route for practical reasons: we and others have found that end-to-end VQ training at very low rates can be brittle and resource-intensive [18, 7], and using pre-quantizer latents from a discrete model can complicate diffusion training [27]. Adding RVQ post-hoc leverages the stability of the continuous model and allows the same architecture to encode into and decode from either KL or VQ latents, providing flexibility for downstream applications and facilitating controlled comparisons between diffusion/flow models and language models.
@@ 3.3 Evaluation metrics
We measure processing speed via real-time factor (RTF) and reconstruction quality using scale-invariant signal-to-distortion-ratio (SI-SDR), multi-resolution short-time Fourier transform (STFT) loss, mel-spectrogram L1 distance, PESQ-WB, and Audiobox Aesthetics, all in bfloat16 precision. We choose bfloat16 over float32 or float64 as nearly all generative modeling is performed in reduced precision, and is our target application. Note that float32 yields significantly different values but similar trends. For baselines at non-44.1 kHz rates, we resample to native rates for processing and back to 44.1 kHz for evaluation. We use the Song Describer [16] dataset and filter out all vocals using a voice activity detection model. We only train on instrumental content and aim to only evaluate on instrumental content.
@@ 3.4 Baseline comparisons
We evaluate GenAE through speed, audio channel format, and latent format ablation studies. We compare with a suite of state-of-the-art codecs. Baselines include DAC, EnCodec (24/48 kHz), Stable Audio Open, CoDiCodec (discrete/continuous), MOSS, and UniAudio2.
@@ 4.1 Speed benchmark
Our generative-first design philosophy prioritizes encoding speed, recognizing that generative models require encoding massive datasets during training. This makes encoding speed key to training throughput and to on-the-fly data augmentations (e.g. stem remixing, pitch/time shifts) where precomputing all possible augmentation-combinations would be impractical. We evaluate design decisions using real-time factor (processed/elapsed time). Fig. 1 shows results for encoding and decoding 2 60-second stereo samples at 44.1 kHz. Starting from our DAC-like baseline with Snake activations (Base), we achieve progressive improvements through: efficient activations (+5% encode), pre-residual downsampling (+36% encode), separable convolutions (+7% encode), and aggressive temporal downsampling (15 encoding speedup). Additional components include mel-spectrogram fusion and windowed attention with format conditioning. The resulting GenAE model uses 3 less memory than SAO, encodes over 10 faster than SAO, and decodes 1.6 faster than DAC, enabling faster processing for easier generative modeling.
@@ 4.2 Compression and quality benchmark
Effective generative modeling requires autoencoders that achieve aggressive compression and good reconstruction quality while supporting continuous and discrete latents, enabling diffusion or language modeling. Our evaluation demonstrates how GenAE achieves this joint objective. Fig. 3 shows GenAE achieving superior rate-distortion performance essential for generative modeling. At 13.125 Hz, GenAE surpasses SAO in all metrics while running at 60% of the rate, and matches DAC performance at just 15% of the rate. This directly enables faster training and reduced memory requirements for generative models. At 36.75 Hz, GenAE-KL outperforms all baselines across all metrics while maintaining substantial compression. GenAE-VQ at 13.125 Hz again outperforms SAO across all metrics, demonstrating unified architecture performance across both continuous and discrete representations. Note that CoDiCodec sounds better than the metrics show. This unification eliminates the need for separate models when comparing diffusion and language model approaches, addressing a benchmarking challenge [22]. We evaluate PESQ-WB at 16 kHz following [14]. GenAE-KL maintains its rate-distortion frontier position, with 36.75 Hz achieving the best score (4.04) and 13.125 Hz providing strong high-compression performance (3.00). Baselines score: DAC (3.49), EnCodec-48 (3.77), EnCodec-24 (3.67), SAO (2.76), MOSS (2.82), CoDiCodec (1.73), UniAudio2 (1.31). We conduct reference-free evaluation using the Audiobox Aesthetics model [24], which reveals tighter clustering. GenAE-KL scores 7.87 (36.75 Hz) and 7.88 (13.125 Hz), while baselines range from 7.77-7.88: MOSS (7.88), EnCodec-48 (7.83), SAO (7.82), EnCodec-24 (7.80), and CoDiCodec/DAC/UniAudio2 (7.77-7.78). Both metrics operate at 16 kHz, which may favor lower-bandwidth models like EnCodec-24, MOSS, and UniAudio2.
@@ 4.3 Multi-format unification benchmark
Practical generative modeling requires autoencoders that work seamlessly across different audio channel configurations. Our unified architecture eliminates the need for multiple specialized models, reducing deployment complexity with consistent performance across major audio channel formats. Tab. 1 evaluates mel-spectrogram reconstruction performance across L/R and M/S stereo audio channel formats. GenAE performs nearly identically in both -KL and -VQ forms across audio channel formats, demonstrating consistent quality regardless of input audio channel format or internal latent configuration. This is crucial for generative models to process diverse content. In contrast, CoDiCodec, SAO and EnCodec-48 degrade significantly due to L/R-specific training, while mono models are varied (DAC/UniAudio2/MOSS maintain, EnCodec-24 degrades). This highlights the importance of multi-format design since GenAE-based models can manipulate different audio formats without needing to re-encode/decode.
@@ 4.4 Generative context benchmark
| EnCodec-48 | 150 | 73 | 0.5485 | 0.6601 |
| DAC | 86 | 127 | 0.5114 | 0.5144 |
| CoDiCodec-FSQ | 11 | 993 | 0.9596 | 1.0558 |
| MOSS | 12.5 | 873 | 0.6945 | 0.6958 |
| UniAudio2 | 12.5 | 873 | 2.0371 | 2.0561 |
@@ 5 Conclusion
In this work, we design neural audio autoencoders for generative models. Our GenAE is a single backbone that autoencodes at rates as low as 13.125 Hz ( downsampling), supports both continuous (KL) and discrete (VQ) latents, and operates across mono, stereo, and mid/side audio channel formats without model variants. A principled combination of efficient activations, early downsampling, separable convolutions, mel-spectrogram fusion, and windowed attention yields order-of-magnitude faster encoding and substantially lower memory while maintaining competitive fidelity, placing GenAE on the rate-distortion frontier with far fewer tokens. Our generative-first design repositions tokenization from a computational bottleneck to an efficient, scalable component: a 60-second track compresses to 788 tokens, enabling tractable long-context transformers. A post-training step lets us use both continuous and discrete latents in one model.